# HJRNO: Hamilton-Jacobi Reachability with Neural Operators

### v1

@@ Abstract
Ensuring the safety of autonomous systems under uncertainty is a critical challenge. Hamilton-Jacobi reachability (HJR) analysis is a widely used method for guaranteeing safety under worst-case disturbances. Traditional HJR methods provide safety guarantees but suffer from the curse of dimensionality, limiting their scalability to high-dimensional systems or varying environmental conditions. In this work, we propose HJRNO, a neural operator-based framework for solving backward reachable tubes (BRTs) efficiently and accurately. By leveraging the Fourier Neural Operator (FNO), HJRNO learns a mapping between value functions, enabling fast inference with strong generalization across different obstacle shapes, system configurations, and hyperparameters. We demonstrate that HJRNO achieves low error on random obstacle scenarios and generalizes effectively across varying system dynamics. These results suggest that HJRNO offers a promising foundation model approach for scalable, real-time safety analysis in autonomous systems.
@@ 1 Introduction
The computation of the BRT involves solving the Hamilton-Jacobi (HJ) partial differential equation (PDE) over the state space, typically discretized using a grid-based dynamic programming approach. This traditional method can produce highly accurate results when a high-resolution grid is used. However, the “curse of dimensionality” poses a significant challenge, as the number of grid points grows exponentially with the dimension of the system. Moreover, solving the HJ-PDE requires cross-dimensional computations, further compounding the computational burden.
@@
Recently, deep learning-based approaches have been proposed to efficiently compute the BRT. The physics-informed machine learning framework [2, 3, 13, 14] introduces neural PDE solvers that leverage neural networks to compute the BRT. However, these frameworks only learn the solution function for a single instance rather than the underlying functional mapping. Consequently, when the input function changes, a new neural PDE solver must be trained from scratch.
@@
    By learning a neural operator, we achieve near-instantaneous inference ( $10^{-3}$ seconds), eliminating the need to solve the HJ-PDE for each different problem setting, such as varying obstacle shapes or system hyperparameters.
@@
    We demonstrate that our method generalizes well across randomly generated obstacle shapes and varying system hyperparameters, indicating strong potential for applying neural operators to BRT problems.
@@ 2 Problem Setting
where $x$ denotes the system state, $u$ is the control input, and $d$ represents an external disturbance. The sets $\mathcal{U}$ and $\mathcal{D}$ denote the sets of measurable control and disturbance functions, respectively. We define the system trajectory starting from initial state $x$, subject to control $u(\cdot)$ and disturbance $d(\cdot)$ over the time interval $[t,\tau]$, as $\zeta(x,u,d,t,\tau)$. The unsafe set is defined as =$\mathcal{L}=\{x\;:\;l(x)\leq 0\}$. The backward reachable tube (BRT) is the set of all initial states from which, despite applying optimal control strategies aimed at avoiding $\mathcal{L}$, there exists a disturbance strategy that forces the system into the unsafe set within the time horizon $[t,\tau]$:
$\mathcal{R}(t)=\{x:\forall u(\cdot),\exists d(\cdot),\exists\tau\in[t,T],\zeta(x,u,d,t,\tau)\in\mathcal{L}\}$ (2)
Since the backward reachable tube (BRT) contains unsafe states, the objective in system design is to steer trajectories away from the BRT. To compute the BRT, we first define an objective function that measures the minimum value of $l$ along a trajectory:
$J(x,t,u,d)=\underset{\tau\in[t,T]}{\mathrm{min}}l(\zeta(x,u,d,t,\tau))$ (3)
[…]
$V(x,T)$ $=l(x)$ (4b)
$\min\left\{D_{t}V(x,t)+H(x,t),\;l(x)-V(x,t)\right\}=0$ (5)
where the Hamiltonian $H(x,t)$ is defined as:
[…]
As an illustrative example, we model an autonomous vehicle using a generalized Dubins car model. Unlike the classical Dubins car, this formulation allows the vehicle’s speed to vary and adopts a more realistic heading control mechanism based on curvature, which is standard for nonholonomic vehicles. Specifically, rather than directly controlling the turning rate, we apply a curvature control scaled by the velocity. In the following, we refer to this Dubins-like vehicle with acceleration and curvature control as the dynamic Dubins car. The system dynamics are given by:
@@
$l(x)=(x_{1}-x_{\mathrm{tree}})^{2}+(x_{2}-y_{\mathrm{tree}})^{2}-r$ (8)
@@ 3 Method
3.1 Neural Operator
Operator learning focuses on learning mappings between coefficient functions (inputs) and solution functions (outputs). A straightforward approach would be to discretize these functions onto grids, allowing the use of conventional neural network architectures such as multilayer perceptrons (MLPs), convolutional neural networks (CNNs), or vision transformers (ViTs) for learning. However, this strategy inherently ties the model to the specific discretization used during training, limiting its ability to generalize across different resolutions or grid structures.
In contrast, neural operators are designed to directly learn mappings between function spaces without relying on fixed discretizations. As a result, neural operators can naturally adapt to new discretizations and varying resolutions without the need for retraining. Moreover, traditional neural networks often require significant architectural modifications to maintain a consistent level of accuracy across different discretizations. Neural operators are the only known models that offer a theoretical guarantee of both universal approximation and invariance to discretization [9].
The Fourier transform simplifies differentiation by converting it into multiplication in the Fourier domain, making it a popular tool for solving PDEs. Building on this principle, the Fourier Neural Operator (FNO) [11] has emerged as one of the leading architectures for learning operators between function spaces. The FNO constructs the mapping using a sequence of $L$ integral kernel operator layers:
$s=\mathcal{M}_{\theta}(a):=\mathcal{Q}\circ\sigma(\mathcal{W}_{L}+\mathcal{K}_{L})\circ\cdots\circ\sigma(\mathcal{W}_{1}+\mathcal{K}_{1})\circ\mathcal{P}(a)$ (9)
where $a(\cdot)$ and $s(\cdot)$ are the input and output functions, $\sigma$ is an activation function, $\mathcal{P}$ and $\mathcal{Q}$ are shallow fully connected networks for projecting to and from higher-dimensional spaces, and $\mathcal{W}_{i}$ are learned linear maps. Each $\mathcal{K}_{i}$ represents a Fourier integral operator to the intermediate layer output $v_{i}$:
$(\mathcal{K}_{i}v_{i})(x)=\mathcal{F}^{-1}\big{(}R_{i}\cdot(\mathcal{F}v_{i})\big{)}(x)$ (10)
where $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the Fourier and inverse Fourier transforms, respectively, and $R_{i}$ is a learnable weight matrix that parameterizes the kernel function in the Fourier space. By utilizing frequency-domain representations, FNO efficiently capture both local and global behaviors of complex functions.
3.2 HJRNO
We apply the Fourier Neural Operator (FNO) [11] to learn the functional map for backward reachable tubes (BRT):
$a(x)$ $:=V(x,\,t=T)$ (11a)
$s(x)$ $:=V(x,\,t=0)$ (11b)
To construct the training dataset, we first compute the solutions using traditional dynamic programming solvers. We then train the FNO (9) in a purely data-driven manner to learn the mapping between these functions. Although intermediate time steps are not explicitly available, we set $T$ sufficiently large so that $V(x,\,t=0)$ stabilizes and does not change with further increases in $T$. This setup reflects the most common practical use case.
Once trained, our method, which we call HJRNO (Hamilton-Jacobi Reachability with Neural Operators), achieves an inference time of approximately $\sim 10^{-3}$ seconds—orders of magnitude faster than traditional solvers, which typically require several seconds to hours per evaluation.
Additionally, to account for hyperparameters $h$ (e.g., the maximum acceleration limit in the dynamic Dubins car model from the Problem Setting section), we extend the input by appending $h$ as a constant feature across all spatial locations:
$a(x)$ $:=V(x,\,t=T,h)$ (12a)
$s(x)$ $:=V(x,\,t=0,h)$ (12b)
Leveraging the universal approximation and discretization invariance properties of FNO, HJRNO can predict solutions for a wide range of $V(x,\,t=T,h)$ instances. As we show in the Results section, HJRNO exhibits strong interpolation capability even with a limited training set, greatly reducing computational costs when handling varying obstacle shapes or system hyperparameters—scenarios where traditional BRT recomputation would be expensive. Finally, thanks to discretization invariance, HJRNO also enables zero-shot super-resolution: it can predict high-resolution solutions even when trained only on low-resolution datasets.
@@ 4 Results
Ground TruthRefer to captionRefer to captionRefer to captionPredictionRefer to captionRefer to captionRefer to captionError MapRefer to captionRefer to captionRefer to caption
Figure 1: Predicted BRTs for randomly selected test cases from the random obstacle shape experiment
We evaluate HJRNO on the dynamic Dubins car model introduced in the Problem Setting section. To simulate realistic and diverse environments, we generate random obstacles with smooth, arbitrary shapes, examples of which are shown in Fig.4. The model is trained on 100 such obstacle shapes. Across 1000 test cases, HJRNO achieves a low average error of 0.015. Results randomly selected from the test dataset are presented in Fig.1.
We further evaluate HJRNO’s ability to generalize across varying system hyperparameters, specifically the maximum acceleration limit and maximum curvature limit. In this experiment, we use a fixed circular obstacle shape. We sample a sparse training dataset by selecting 10 uniformly spaced values along each hyperparameter axis, yielding $10\times 10=100$ training points. The range of each hyperparameter is shown in Fig. 3.

### v2

@@ Abstract
Ensuring the safety of autonomous systems under uncertainty is a critical challenge. Hamilton-Jacobi reachability (HJR) analysis is a widely used method for guaranteeing safety under worst-case disturbances. In this work, we propose HJRNO, a neural operator-based framework for solving backward reachable tubes (BRTs) efficiently and accurately. By leveraging neural operators, HJRNO learns a mapping between value functions, enabling fast inference with strong generalization across different obstacle shapes and system configurations. We demonstrate that HJRNO achieves low error on random obstacle scenarios and generalizes effectively across varying system dynamics. These results suggest that HJRNO offers a promising foundation model approach for scalable, real-time safety analysis in autonomous systems.
footnotetext: This work was supported by the Canada CIFAR AI Chairs and NSERC Discovery Grants Programs.
@@ 1 Introduction
The computation of the BRT involves solving the Hamilton-Jacobi-Isaacs (HJI) partial differential equation (PDE) over the state space, typically discretized using a grid-based dynamic programming approach. This traditional method can produce highly accurate results when a high-resolution grid is used. However, the “curse of dimensionality” poses a significant challenge, as the number of grid points grows exponentially with the dimension of the system. Moreover, solving the HJI-PDE requires cross-dimensional computations, further compounding the computational burden.
@@
Recently, deep learning-based approaches have been proposed to efficiently compute the BRT. The physics-informed machine learning framework [3, 23, 24] introduces neural PDE solvers that leverage neural networks to compute the BRT. However, these approaches typically learn the solution function for a single problem instance and must be retrained from scratch when the problem setting changes. While [4] introduces a parameterized neural PDE solver, it only handles scalar hyperparameters rather than functions, limiting its applicability to more general settings where the input includes functions.
@@
    By learning a neural operator, we achieve near-instantaneous inference ( $10^{-3}$ seconds), eliminating the need to solve the HJI-PDE for each different problem setting, such as varying obstacle shapes or system hyperparameters.
@@
    We demonstrate that our method generalizes well across randomly generated obstacle shapes and varying system hyperparameters, indicating strong potential of neural operators for broad applicability in BRT problems.
@@ 2 Problem Setting
where $x$ denotes the system state, $u$ the control input, and $d$ the external disturbance. The sets $\mathcal{U}$ and $\mathcal{D}$ denote the sets of measurable control and disturbance functions, respectively. We define the system trajectory starting from initial state $x$, subject to control $u(\cdot)$ and disturbance $d(\cdot)$ over the time interval $[t,\tau]$, as $\zeta(x,u(\cdot),d(\cdot),t,\tau)$. The unsafe set is defined as
=$\mathcal{L}=\{x\;:\;l(x)\leq 0\},$ (2)
where the function $l(x):\mathbb{R}^{n}\rightarrow\mathbb{R}$ typically denotes the signed distance function. The backward reachable tube (BRT) is the set of all initial states from which, despite applying optimal control strategies aimed at avoiding $\mathcal{L}$, there exists a disturbance strategy that forces the system into the unsafe set within the time horizon $[t,\tau]$:
$\mathcal{R}(t)=\{x:\forall u(\cdot),\exists d(\cdot),\exists\tau\in[t,T],\zeta(x,u(\cdot),d(\cdot),t,\tau)\in\mathcal{L}\}$ (3)
2.1 Running examples
Air3D. This example models a two-vehicle collision avoidance scenario. The system consists of two agents: an evader, which attempts to avoid capture, and a pursuer, which tries to intercept the evader.
The dynamics of the evader (agent $a$) and pursuer (agent $b$) are given by:
$\begin{bmatrix}\dot{x}_{a}\\ \dot{y}_{a}\\ \dot{\theta}_{a}\end{bmatrix}=\begin{bmatrix}v\cos(\theta_{a})\\ v\sin(\theta_{a})\\ u_{a}\end{bmatrix},\quad\begin{bmatrix}\dot{x}_{b}\\ \dot{y}_{b}\\ \dot{\theta}_{b}\end{bmatrix}=\begin{bmatrix}v\cos(\theta_{b})\\ v\sin(\theta_{b})\\ u_{b}\end{bmatrix}$ (4)
Here, $v$ is the constant speed shared by both agents, while $u_{a}$ and $u_{b}$ are the angular velocities of the evader and pursuer, respectively, with control bounds $u_{a},\;u_{b}\in[-u_{\mathrm{max}},\,u_{\mathrm{max}}]$.
To simplify the problem, we can formulate the system dynamics in the evader’s frame of reference, yielding a relative dynamics model:
$\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\\ \dot{x}_{3}\end{bmatrix}=\begin{bmatrix}-v_{a}+v_{b}\cos(x_{3})+u_{a}x_{2}\\ v_{b}\sin(x_{3})-u_{a}x_{1}\\ u_{b}-u_{a}\end{bmatrix}$ (5)
In this relative system, $x_{1}$ and $x_{2}$ represent the position of the pursuer relative to the evader, and $x_{3}$ is the relative heading angle.
The unsafe set is defined as:
=$\mathcal{L}=\left\{(x_{1},x_{2})\in\mathbb{R}^{2}\;\middle|\;\|(x_{1},x_{2})\|\leq d\right\}$ (6)
which corresponds to the set of states where the evader and pursuer are within a distance $d$ of each other, i.e. a collision.
The backward reachable tube (BRT) computed from this unsafe set (3) represents the set of initial states from which the pursuer can guarantee a collision with the evader, despite the evader applying optimal control to avoid it.
Dynamic Dubins car. We model an autonomous vehicle using a generalized Dubins car model. Unlike the classical Dubins car, this formulation allows the vehicle’s speed to vary and adopts a more realistic heading control mechanism based on curvature, which is standard for nonholonomic vehicles. Specifically, rather than directly controlling the turning rate, we apply a curvature control scaled by the velocity. In the following, we refer to this Dubins-like vehicle with acceleration and curvature control as the dynamic Dubins car. The system dynamics are given by:
@@ 2.1 Running examples
=$\mathcal{L}=\left\{(x_{1},x_{2})\in\mathbb{R}^{2}\;\middle|\;\|(x_{1}-x_{\mathrm{tree}},\,x_{2}-y_{\mathrm{tree}})\|\leq r\right\}$ (8)
@@
3 Background
3.1 Hamilton-Jacobi Reachability
To compute the BRT in (3), we first define an objective function that measures the minimum value of $l$ (see Eq. (2) for its definition) along a trajectory:
$J(t,x,u(\cdot),d(\cdot))=\underset{\tau\in[t,T]}{\mathrm{min}}l(\zeta(t,\tau,x,u(\cdot),d(\cdot)))$ (9)
$V(t,x)$ $=\underset{d(\cdot)\in\mathbb{D}}{\mathrm{inf}}\,\underset{u(\cdot)\in\mathbb{U}}{\mathrm{sup}}\;J\left(t,x,u(\cdot),d(\cdot)\right)$ (10a)
$V(T,x)$ $=l(x)$ (10b)
$\min\left\{\frac{\partial}{\partial t}V(t,x)+H(t,x),\;l(x)-V(t,x)\right\}=0$ (11)
where the Hamiltonian $H(t,x)$ is defined as
$H(t,x)=\underset{u(\cdot)\in\mathbb{U}}{\mathrm{max}}\,\underset{d(\cdot)\in\mathbb{D}}{\mathrm{min}}\;\left(\frac{\partial}{\partial x}V(t,x)\right)^{\top}f(x,u(\cdot),d(\cdot))$ (12)
3.2 Neural Operators
Operator learning aims to learn mappings between input function and output function. A naive approach would be to discretize these functions onto grids, and then learn a function approximator – like conventional neural network architectures such as multilayer perceptrons (MLPs), convolutional neural networks (CNNs), or vision transformers (ViTs) – that maps the discretized inputs to the discretized outputs. However, this strategy inherently ties the model to the specific discretization used during training, limiting its ability to generalize across different resolutions or grid structures.
Neural operators are designed to directly learn mappings between function spaces without relying on fixed discretizations. As a result, neural operators can naturally adapt to new discretizations and varying resolutions without the need for retraining.
Neural operators are the only known models that offer a theoretical guarantee of both universal approximation and invariance to discretization [16]. In this context, universal approximation refers to the ability of neural operators to approximate any continuous operator between Banach spaces to arbitrary accuracy, while discretization invariance ensures the model generalizes across different resolutions without retraining.
Formally, the objective of neural operators is to learn a mapping between input and output functions defined over infinite-dimensional spaces. Let $a\in\mathcal{A}$ and $s\in\mathcal{S}$ denote the input and output functions, respectively, where $\mathcal{A}$ and $\mathcal{S}$ are Banach spaces. The true underlying operator is denoted by $\mathcal{M}^{\ast}$, such that
$s(\cdot)=\mathcal{M}^{\ast}(a(\cdot))$ (13)
The goal is to approximate this operator using a parameterized model $\mathcal{M}_{\theta}:\;\mathcal{A}\rightarrow\mathcal{S}$, where $\theta\in\mathbb{R}^{P}$ denotes the learnable parameters. This approximation is obtained by minimizing the empirical risk over a dataset of $N$ input-output function pairs:
[…]
Several neural operator architectures have been proposed to approximate mappings between infinite-dimensional function spaces [20, 19, 14, 11, 1, 16]. In this work, we focus on two prominent approaches: Transformer-based Neural Operator (TNO) [7] and the Fourier Neural Operator (FNO) [18]. TNO model complex spatial correlations via attention mechanisms, while FNOs use Fourier transforms to capture global structure in the spectral domain.
3.2.1 Fourier Neural Operator
The FNO constructs the mapping using a sequence of $L$ integral kernel operator blocks:
$s=\mathcal{M}_{\theta}(a):=\mathcal{P}_{\text{out}}\circ\text{Block}_{L}\circ\cdots\circ\text{Block}_{1}\circ\mathcal{P}_{\text{in}}(a)$ (15)
where $a(\cdot)$ and $s(\cdot)$ are the input and output functions, $\mathcal{P}_{\text{in}}$ and $\mathcal{P}_{\text{out}}$ are shallow fully connected networks for projecting to and from higher-dimensional spaces.
Each block takes the form:
$\text{Block}_{i}(f)(x):=\sigma\left(W_{i}\cdot f(x)+\mathcal{K}_{i}(f)(x)\right)$ (16)
where $\sigma$ is an activation function, $W_{i}$ is a learnable linear map, and $\mathcal{K}_{i}$ is an integral operator:
[…]
By the convolution theorem [6], this operation is equivalent to:
$\mathcal{K}_{i}(f)(x)=\mathcal{F}^{-1}\big{(}\mathcal{F}(\kappa_{\theta_{i}})\cdot\mathcal{F}(f)\big{)}(x)$ (18)
where $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the Fourier and inverse Fourier transforms, respectively. FNO parameterizes $\mathcal{F}(\kappa_{\theta_{i}})$ using a learnable weight tensor $R_{\theta_{i}}\in\mathbb{R}^{d\times d}$, enabling global convolution in the spectral domain.
3.2.2 Transformer-based Neural Operator
The Transformer-based Neural Operator (TNO) construct mappings between function spaces using a sequence of attention based blocks.
We adopt the Galerkin Transformer architecture [7] as our implementation of the TNO in this work. The neural operator is defined as a composition of projection layers and attention-based blocks:
$s=\mathcal{M}_{\theta}(a):=\mathcal{P}_{\text{out}}\circ\text{Block}_{L}\circ\cdots\circ\text{Block}_{1}\circ\mathcal{P}_{\text{in}}(a)$ (19)
Each block is constructed as:
$\text{Block}_{i}(f):=f+\text{Attn}_{i}(f)+\mathcal{P}_{i}(f+\text{Attn}_{i}(f))$ (20)
where $\mathcal{P}_{\text{in}}$ $\mathcal{P}_{\text{out}}$, and $\mathcal{P}_{i}$ are shallow projection networks (MLPs), and $\text{Attn}_{i}$ denotes a self-attention operator at layer $i$.
The attention mechanism approximates an integral operator over the domain $D$ via:
[…]
Here
  1. 1.
    $Q=XW_{Q}$, $K=XW_{K}$, and $V=XW_{V}$ are the query, key, and value matrices. Layer normalization is applied to $K$ and $V$.
  2. 2.
    $X\in\mathbb{R}^{n\times d}$ is the input tensor, representing the function sampled at $n$ spatial points with $d$ dimensional features.
  3. 3.
    $W_{Q}$, $W_{K}$, $W_{V}\in\mathbb{R}^{d\times d}$ are learnable projection weights.
  4. 4.
    $K_{i}$ denotes the $i$th feature column of $K$.
  5. 5.
    $Q_{ij}$ denotes the attention weight corresponding to the $i$th feature and the $j$th spatial location. In Eq. (21), we interpret the spatial location indexed by $j$ as the point $x$ where the attention output is being evaluated.
@@ 4 Method
We use the Fourier Neural Operator (FNO) [18] and Transformer-based Neural Operator (TNO) [7] to learn a functional mapping for backward reachable tubes (BRTs). We define the initial time as $t=0$ and the final time as $t=-T$. We assume that the value function $V(t,x)$, which satisfies the HJI-PDE (11), converges as $T\rightarrow\infty$. We denote this limiting solution as
$V_{\infty}(x):=\lim_{T\rightarrow\infty}V(t=-T,x),$ (22)
which corresponds to the maximal BRT, and is most commonly used in practice for providing safety filters [15, 5].
Our goal is to learn an operator $\mathcal{M}_{\theta}$ that maps the initial value function $V(t=0,x)$ to the solution $V_{\infty}(x)$:
$a(x)$ $:=V(t=0,\,x)$ (23a)
$s(x)$ $:=V_{\infty}(x)$ (23b)
To construct the training dataset $\mathcal{D}=\left\{\left(V^{(i)}(t=0,x),\,V_{\infty}^{(i)}(x)\right)\right\}_{i=1}^{N}$, we solve the HJI-PDE in Eq. (11) for various problem instances using traditional dynamic programming solvers [21]. The Neural Operator model (15) (19) is then trained in a data-driven manner to approximate the mapping $\mathcal{M}_{\theta}:\;V(t=0,\,x)\mapsto V_{\infty}(x)$.
Both $V(t=0,\,x)$ and $V_{\infty}(x)$ are represented by uniformly sampling the value function over a bounded domain $D\subset\mathbb{R}^{n}$ in the state space of $x$. For each training function instance $i$, the function is represented as a set of values on a uniform grid:
$\left\{V^{(i)}(t=0,x_{1}),V^{(i)}(t=0,x_{2}),\dots V^{(i)}(t=0,x_{M})\right\}$ (24)
where $\{x_{j}\}^{M}_{j=1}$ are grid points in the domain $D$. The same sampling is applied to represent $V^{(i)}_{\infty}(x)$.
The model is trained by minimizing the mean squared error between predicted and ground truth value functions:
[…]
4.1 Generalizing to Parametric Inputs
Our method is able to handle scalar hyperparameters by embedding them as constant input functions over the state domain. That is, for a given hyperparameter $h\in\mathbb{R}^{m}$ (e.g., the maximum acceleration limit in the dynamic Dubins car example from Section Running examples), we define an input function $\tilde{h}(x)=h$ for all $x\in D$. This allows our framework to incorporate parametric dependence using the same neural operator architecture.
In this setup, we train the model to learn the mapping:
$a(x)$ $:=V(t=0,\,x,\,h),$ (26a)
$s(x)$ $:=V_{\infty}(x,\,h),$ (26b)
where the value function now depends on both the state $x$ and the hyperparameter $h$.
@@ 5 Results
We evaluate HJRNO on 6 distinct experimental setups, each designed to test generalization under different conditions such as random geometry and system dynamics. In all experiments, the model takes the initial value function as input and predicts the value function at the infinite time horizon (see Eq. (23)). This function typically varies only with the spatial coordinates $(x_{1},x_{2})$, while remaining approximately constant along other dimensions such as velocity or heading.
To improve efficiency, we treat the spatial coordinates $(x_{1},x_{2})$ as the primary input domain and encode the remaining state variables as constant hyperparameter $h\in\mathbb{R}^{m}$, following the formulation in Eq. (26). This approach significantly improves both performance and scalability for FNO and TNO models. In the Air3D experiment, where the full state space includes position $(x,y)$ and heading $\theta$, using the full state led to notable degradation: FNO’s test error increased from 0.028 to 0.050, training time rose from 1 minute to 3 minutes, and model size grew from 38 MB to 454 MB; Similarly, TNO’s test error increased from 0.037 to 0.15, with training time rising from 4 to 6 minutes.
All experiments are conducted at a resolution of 50, with each model trained for 20 epochs. A summary of quantitative results across all setups is provided in Table 1 and Table 2. Our method demonstrates significant efficiency: the FNO checkpoint size is roughly equal to the memory required to store one single solution instance of the 4D Dubins car ($50\times 50\times 50\times 50$ array, approx. 25MB), and the TNO model is even smaller. Both FNO and TNO perform inference three orders of magnitude faster than traditional solvers. This efficiency is particularly valuable in robotics applications, where onboard devices often have limited memory and require rapid, high-frequency updates as the environment evolves, such as when obstacles move or change shape.
Table 1: Performance metrics. Train Err and Test Err report relative $L_{2}$ errors on the training and test datasets, respectively. Dataset Time refers to the time required to generate both training and testing datasets using traditional solvers. Train Time denotes the total training duration for 20 epochs. Params indicates the number of trainable parameters in each model.
Method Experiment Train Err Test Err Dataset Time Train Time Params FNO Air3D 0.017 0.028 50 min 1 min 4.83M Single Obstacle 0.0019 0.0092 3 min 17 min 4.83M Two Obstacles 0.0020 0.014 3 min 17 min 4.83M Indoor Environment 0.0035 0.029 20 min 40 min 4.83M Velocity-Dependent 0.0021 0.035 11 min 40 min 4.83M Parametric Inputs 0.0007 0.016 3 min 20 min 4.83M TNO Air3D 0.034 0.037 50 min 4 min 0.8M Single Obstacle 0.0014 0.0020 3 min 80 min 0.8M Two Obstacles 0.0024 0.0063 3 min 80 min 0.8M Indoor Environment 0.0062 0.023 20 min 190 min 0.8M Velocity-Dependent 0.0025 0.044 11 min 187 min 0.8M Parametric Inputs 0.0004 0.013 3 min 95 min 0.8M
Table 2: Resource usage metrics. VRAM (Train) denotes the peak GPU memory usage during training. Data Size refers to the storage size of the training dataset. Chkpt Size is the size of the saved model checkpoint. Solve Time is the time required to compute a single solution using traditional HJR solvers. Inference Time is the time required for the trained model to produce a prediction.
Method Experiment VRAM (Train) Data Size Chkpt Size Solve Time Inference Time FNO Air3D 0.5 GB 49 MB 38 MB 0.05 s 0.001 s Single Obstacle 0.5 GB 2.4 GB 38 MB 0.9 s 0.003 s Two Obstacles 0.5 GB 2.4 GB 38 MB 1.0 s 0.002 s Indoor Environment 0.5 GB 7.3 GB 38 MB 2.6 s 0.003 s Velocity-Dependent 0.5 GB 7.3 GB 38 MB 1.0 s 0.002 s Parametric Inputs 0.5 GB 2.4 GB 38 MB 0.9 s 0.002 s TNO Air3D 1.8 GB 49 MB 3 MB 0.05 s 0.003s Single Obstacle 1.8 GB 2.4 GB 3 MB 0.9 s 0.004 s Two Obstacles 1.8 GB 2.4 GB 3 MB 1.0 s 0.003 s Indoor Environment 1.8 GB 7.3 GB 3 MB 2.6 s 0.005 s Velocity-Dependent 1.8 GB 7.3 GB 3 MB 1.0 s 0.004 s Parametric Inputs 1.8 GB 2.4 GB 3 MB 0.9 s 0.005 s
5.1 Air3D Collision Avoidance
The Air3D problem setup is detailed in Section Running examples. In this experiment, we test HJRNO’s ability to generalize over varying agent geometries. Specifically, we randomly generate arbitrary smooth shapes for both the evader and the pursuer, as visualized in Fig. 1.
Each shape is created by sampling random radii at evenly spaced angles in polar coordinates, applying a convex hull to ensure a realistic boundary, and then smoothing the resulting polygon using cubic B-spline interpolation.
Using these shapes, we generate 100 random pairs of evader and pursuer geometries. The training dataset consists of the corresponding $(V(t=0,x),V_{\infty}(x))$ pairs. The model is trained on these 100 samples and evaluated on 50 additional, unseen pairs. Examples of results are shown in Fig. 8.
5.2 Dynamic Dubins car
All five experiments in this section are based on the dynamics Dubins car model described in the Running examples section.
Single Obstacle.
We train on 100 randomly generated smooth obstacles and test on 50. Obstacles are created using the same random shape generation method as in the Air3D experiment (see Fig.1). Examples of results are shown in Fig.2.
Two Obstacles.
This experiment introduces a second obstacle with a randomly sampled separation distance from the first. A visualization is shown in Fig. 6. Note that simply taking the union of the solutions for each obstacle individually does not yield the correct solution for the combined two-obstacle scenario, as shown by Fig. 4. We generate 100 training and 50 testing samples of $(V(t=0,x),V_{\infty}(x))$ pairs.
Indoor Environment.
We simulate indoor spaces featuring walls, doors, and interior clutter. Room layouts are generated with randomized door placements and walls, and obstacles inside the room include randomly sized boxes, circles, ellipses. We generate 300 training and 50 testing $(V(t=0,x),V_{\infty}(x))$ pairs. Examples of results are shown in Fig. 10. Fig. 9 illustrates that neural operators can perform zero-shot super-resolution, whereas conventional networks (e.g., CNNs) fail to generalize across resolutions.
Velocity-Dependent Obstacles.
In this setup, the obstacle size increases as a function of velocity, simulating environments where faster motion demands greater clearance. For each sample, the obstacle radius grows with velocity according to either an exponential or logarithmic function, with both the function type and parameters randomly sampled. An illustration is shown in Fig.7. This setup introduces velocity-dependence in the value function. We train on 300 and test on 50 such $(V(t=0,x),V_{\infty}(x))$ pairs. Examples of predictions are provided in Fig. 11.
Parametric Inputs.
To assess generalization over continuous hyperparameter spaces, we vary two control limits: maximum acceleration and maximum curvature. The training set consists of $10\times 10=100$ combinations from uniformly sampled values along each axis. We evaluate on 50 test samples randomly drawn with a uniform distribution along the diagonal of the hyperparameter space. This setting highlights interpolation ability over hyperparameter inputs. See Fig. 3 for the sampling strategy and Fig. 5 for example predictions.
Figure 1: Random obstacle shapes
Ground TruthPredictionError MapFNORefer to captionRefer to captionRefer to captionTNORefer to captionRefer to captionRefer to caption
Figure 2: Single Obstacle
Figure 4: